The rand() function generates a random real number between 0 and 1, but what if we want real numbers in a different range? Use the rand function to generate a 4x5 matrix with random real numbers between 5 and 10. Store your result in the variable "mat".
Answer #1
% Code
mat = (10-5)*rand(4,5)+5

Output:

octave: 1> mat = (10-5) *rand (4,5) +5 mat = 7.1797 6.3239 7.3478 9.4018 9.7492 5.7504 6.5630 9.9115 6.1135 9.3635 8.8120 6.5

Note : Please comment below if you have any doubts, upvote the solution if it helped. Thanks!